OpenStudy (anonymous):

Betty is having invitations printed for a holiday party. It costs $4.35 to set up the printing machine and $3.37 per invitation. In the equation below, x represents the number of invitations Betty is having printed, and y represents the cost to print the invitations. y = $3.37x + $4.35 How much will it cost Betty to print 125 invitations? $129.35 $547.12 $429.95 $425.60

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

y=3.37(125)+4.35=?

OpenStudy (anonymous):

Someone answer this please!

OpenStudy (dtan5457):

You are unable to solve y=3.37(125)+4.35=?

OpenStudy (anonymous):

Yes im so sorry see im in a rush thats why

OpenStudy (dtan5457):

That is an equation to type in a calculator and get an answer within 10 seconds.

OpenStudy (dtan5457):

I'm not allowed to give answers.

OpenStudy (anonymous):

Sir dont makee me talk words i shouldnt

OpenStudy (dtan5457):

Looks like all those direct answers have made you lazy, try again.

OpenStudy (anonymous):

Excuse me?

OpenStudy (dtan5457):

Your still on the "easy" side of math, please take advantage of it.

OpenStudy (dtan5457):

By now, you would have solved 10 of these questions if you knew how to do them.

OpenStudy (anonymous):

By now you would of answered them and also if you dont like the way i want it i can care a less just dont come to my question!

OpenStudy (dtan5457):

*I couldn't care less. Bit of a misconception to say "I can care less'

OpenStudy (anonymous):

I can care a less, I can care a less, I can care a less :)

OpenStudy (dtan5457):

That implies that you DO care. Not even the point, however.

OpenStudy (anonymous):

Hun you probably don't know me and you don't want to :) Bye!

OpenStudy (dtan5457):

Your pretty damn right I don't want to. Buh bye. Just saying, I'm reporting every direct answer given ;)

OpenStudy (anonymous):

Report millions I CAN CARE A LESS! :)))))

OpenStudy (dtan5457):

Ok, continue to talk like the smart one here. Continue to fail tests and realize your arrogant mistakes when you reach the "kill myself" level of math. (AKA things 20 times as hard as your questions)

OpenStudy (anonymous):

No not at all i have a 90% in math i needed some help but not from a person like you! I was born with pride and respect and highly educated im supposed to be in 10th grade they switched me to 12th Again i don't care what you think and say about me :) If you dont like it Leave away but i guess you know im right so you cant leave

OpenStudy (dtan5457):

I learned this in 7th grade. Karma strikes back, I promise you.

OpenStudy (anonymous):

Aww i know its going to strike back on you! :) Try hard not to let it.

OpenStudy (dtan5457):

I try to help people, people try to help me, and grateful people thank me for it. Good karma is coming my way.

OpenStudy (anonymous):

Im glad! Bye now :)

OpenStudy (anonymous):

I know ill try not to let it :) !